INSTRUCTIONS Answer all questions on the answer sheet provided You have 2 hours to complete this test *Use of calculators is allowed Standard Normal Distribution table will be provided Q 1. (8 marks) [CLO 1, 2] Ali wants to invest in stocks. He has three alternatives along with different probabilities. The table below shows conditional values for the various alternatives (stocking decision) and states of nature (size of crowd). All values are in US Dollars. Large Average Small Alternative 1 9,000 8,000 1,500 5,000 18,000 12,000 6,000 Alternative 2 15,000 7,000 Alternative 3 If the probabilities associated with the states of nature are 0.20 for a large crowd, 0.30 for an average crowd, and 0.50 for a small crowd, determine: (a) The optimal alternative if the concessionaire is an optimist. (b) The optimal alternative if the concessionaire is a pessimist. (c) The alternative that provides the greatest expected monetary value (EMV). (d) The expected value of perfect information (EVPI). (e) The opportunity loss table. (1 Minimum expected opportunity loss (EOL).
Answer #1

1.

A.

Optimistic decision is the MaxiMax decision.

According to it, maximum of the best payoff under each alternative is chosen.

Maximum of the best payoff under each alternative = $18000 (it is with alternative 3)

So, alternative 3 will be selected under the optimistic criteria.

B.

Pessimist decision is the MiniMax decision.

According to it, minimum of the best payoff under each alternative is chosen.

Minimum of the best payoff under each alternative = $6000 (it is with alternative 3)

So, alternative 3 will be selected under the pessimistic criteria.

C.

EMV for alternative 1= .2*9000 + .3*8000 + .5*(-1500) = $3450

EMV for alternative 2= .2*15000 + .3*7000 + .5*5000 = $7600

EMV for alternative 3= .2*18000 + .3*12000 + .5*6000 = $10200

So, alternative 3 will be selected as it has highest EMV.

D.

EV with PI = .2*18000 + .3*12000 + .5*6000 = $10200

So,

EVPI = WV with PI – EMV = 10200 – 10200 = $0

Pl. repost other unanswered questions for their proper answers!